# Revela: Dense Retriever Learning via Language Modeling

### v1

@@ Abstract
Dense retrievers play a vital role in accessing external and specialized knowledge to augment language models (LMs). Training dense retrievers typically requires annotated query-document pairs, which are costly and hard to obtain in special domains, like code—motivating growing interest in self-supervised retriever learning. Since LMs are trained to capture token-level dependencies through a self-supervised learning objective (i.e., next token prediction), we can analogously cast retrieval as learning dependencies among chunks of tokens. This analogy naturally leads to the question: How can we adapt self-supervised learning objectives in the spirit of language modeling to train retrievers?
To answer this question, we introduce Revela, a unified and scalable training framework for self-supervised retriever learning via language modeling. Revela models semantic dependencies among documents by conditioning next token prediction on both local and cross-document context through an in-batch attention mechanism. This attention is weighted by retriever-computed similarity scores, enabling the retriever to be optimized as part of language modeling. We evaluate Revela on both general-domain (BEIR) and domain-specific (CoIR) benchmarks across various retriever backbones. At a comparable parameter scale, Revela outperforms the previous best method with absolute improvements of 5.2% (18.3% relatively) and 5.6% (14.4% relatively) on NDCG@10, respectively, underscoring its effectiveness. Performance increases with model size, highlighting both the scalability of our approach and its promise for self-supervised retriever learning.
[…]
@@ 1 Introduction
Central to information retrieval are dense retrievers (reimers2019sentence, ; karpukhin-etal-2020-dense, ; ma2024fine, ), which map queries and documents into high-dimensional vector spaces and determine relevance through similarity calculations. Typically, these models rely on carefully annotated query-document pairs and hard negatives for training. However, creating such high-quality training data requires substantial human annotation, which is labor-intensive and difficult to scale in real-world, domain-specific scenarios such as law feng-etal-2024-legal and programming jimenez2024swebench . This limitation stimulates the interest within the community to explore self-supervised approaches for training retrievers directly from unannotated raw texts, i.e., self-supervised retriever learning DBLP:journals/tmlr/IzacardCHRBJG22 .
@@
Modern LMs (grattafiori2024llama, ), a successful case of self-supervised learning, are typically pretrained with the next-token prediction (NTP) paradigm, modeling dependencies among tokens within a single sequence. Analogously, retriever aims to model relationships among larger units—chunks of tokens—capturing more macroscopic dependencies. This motivates us to raise a key question: How can we learn a retriever within a self-supervised learning framework of language modeling? While NTP implicitly identifies the most relevant parts in the context during generation, conditioning language modeling of one sequence on others can serve as an effective proxy for modeling inter-sequence relationships. This can offer a novel and principled approach to self-supervised retriever learning.
@@
In this work, we introduce a self-supervised retriever learning paradigm – Dense Retriever Learning via Language Modeling, abbreviated as Revela.  [^1] Revela merges re- (“retriever”) with vela (“sails” in Latin) to evoke a self-supervised method where retrievers rig the “sails” of language modeling to capture the ambient wind of raw text for label-free learning. As illustrated in Fig.1, Revela trains retrievers by simultaneously optimizing retrievers and LMs; different from conventional NTP, Revela learns the probability of a token given both the prefix in this sequence and all other sequences in the batch. Specifically, in addition to classical self-attention which restricts NTP to individual sequences, an in-batch attention mechanism enables sequences to attend to their neighbors in the same batch during language modeling. In this process, the retriever provides inter-sequence dependencies that modulate the in-batch attention weights, allowing it to be optimized with the LM during training. We split raw texts into chunks within each document and put these chunks into the same batch, motivated by the idea of hard negative samples in contrastive learning (xiong2020approximatenearestneighbornegative, ).
@@
Figure 1: The framework of Revela. Similarity scores between chunks within a batch computed by dense retrievers are used as attention weights for in-batch attention inside Revela’s transformer block. The retriever is trained by optimizing the language modeling objective, i.e., NTP. For instance, when modeling the red sequence, the similarity to the purple chunk will be increased, with related segments highlighted in bold, underline, and italic.
To assess the effectiveness of Revela across diverse retrieval scenarios, we evaluate it on two complementary benchmarks: BEIR (thakur2021beir, ), which covers general-domain retrieval tasks, and CoIR (li2024coir, ), a comprehensive benchmark tailored for code retrieval. Specifically, Revela is trained on Wikipedia for general retrieval and on a set of code-related corpora for code retrieval wang2024coderagbench , using pretrained transformers ranging from 135M to 1B parameters as retriever backbones, with an LM for language modeling. We observe that, with a comparable number of parameters, Revela surpasses REPLUG (shi-etal-2024-replug, ) by an absolute margin of 5.3% on BEIR and 5.4% on CoIR in terms of NDCG@10, corresponding to relative improvements of 18.3% and 14.3%, respectively. These results demonstrate the effectiveness of Revela as a self-supervised retriever learning solution. We also observe that scaling Revela results on the size of retriever backbones and LMs in higher performance gain over baselines. Moreover, Revela maintains consistent performance on both BEIR and CoIR when trained on a mixed-domain dataset. These results collectively highlight Revela’s promising scalability and generalization across both model size and data scale.
To this end, we summarize our contributions as follows:
@@
  • On the BEIR and CoIR benchmarks, Revela surpasses REPLUG, achieving relative NDCG@10 improvements of 18.3% and 14.3%, respectively, underscoring its effectiveness.
  • Revela benefits from both parameter and data scaling: larger retriever and LM backbones yield additional gains, while consistent performance across two benchmarks with mixed-domain training data highlights its generalization.
@@ 2 Related Works
One line of works leverages weak supervision within documents. Contriever (DBLP:journals/tmlr/IzacardCHRBJG22, ) employs contrastive learning, treating passages from the same document as positives and using in-batch negatives, while Co-Condenser (gao-callan-2022-unsupervised, ) warms up Condenser gao-callan-2021-condenser with Wikipedia hyperlinks as pseudo labels. However, these methods are limited by overly simplistic assumptions that structure-associated passages are labeled as pseudo-positives, failing to capture complex relational patterns. Another common strategy is to distill from other retrievers, exemplified by LaPraDoR xu-etal-2022-laprador , which combines dense retrieval with BM25 robertson2009probabilistic through Lexicon-Enhanced Dense Retrieval. Last but not least, autoencoding methods, such as TSDAE (wang-etal-2021-tsdae-using, ) and RetroMAE (xiao-etal-2022-retromae, ), learn sentence embeddings by reconstructing sentences from their neural representations. Autoencoding often lacks pairwise or contrastive supervision, which can result in overfitting to low-level details (steck2020autoencoders, ). Inspired by the aforementioned approaches, our method trains retrievers by modeling dependencies within a document, generating each sequence conditioned on others within the same batch.
@@
LMs (e.g., GPT (NEURIPS2020_1457c0d6, ), LLaMA (touvron2023llama, )) are trained on trillions of tokens with billions of parameters, enabling strong generalization and improving retriever learning through richer semantics and training signals.
LMs can facilitate query and document expansion, as demonstrated by approaches like Query2Doc (wang-etal-2023-query2doc, ) and Doc2Query (gospodinov2023doc2query, ). However, leveraging such capabilities often requires advanced models, leading to high computational costs. Augmenting LMs with relevant information from external knowledge stores not only improves performance in various NLP tasks but also enhances retriever learning. Atlas (izacard2023atlas, ) utilizes cross-attention scores between retrieved documents and the generated output as signal to train the retriever. However, it is based on an encoder-decoder architecture, which diverges from the prevailing trend of decoder-only LMs (touvron2023llama, ; achiam2023gpt, ). Additionally, it requires periodic index updates (reindexing) (guu2020retrieval, ), leading to significant computational costs. In contrast, our work focuses on leveraging the widely adopted decoder-only architecture while eliminating the need for reindexing by modeling relationships between chunks within the same document.
With the rise of decoder-only LMs, REPLUG (shi-etal-2024-replug, ) enhances retrieval by prepending retrieved documents to the queries, training retrievers to produce query-document similarity aligned with the LM’s perplexity. However, the perplexity of frozen LMs is often poorly calibrated (geng-etal-2024-survey, ), resulting in suboptimal retriever learning. This issue can be optimized in Revela where retrievers and LMs are updated jointly during language modeling.
@@
In pre-training corpora, domain-specific knowledge is both scarce and rapidly evolving (grossmann2023ai, ; wen2025synthetic, ), making effective retrieval in specialized domains critically challenging. To enhance the adaptability of dense retrievers across domains, researchers have explored continual learning (sachan-etal-2021-end, ; oguz-etal-2022-domain, ) and task-aware training (cheng-etal-2023-task, ). However, these approaches still rely on query-document pairs from domain-specific datasets. Another approach seeks to simplify domain-specific retrieval for general-purpose dense retrievers. Cai et al. cai-etal-2024-mixgr propose a divide-and-conquer strategy through a mixed-granularity retrieval framework, significantly enhancing dense retriever performance in scientific domains. Our work demonstrates Revela’s domain adaptation capability through language modeling on domain-specific raw texts.
@@ 3.1 Training Objectives
Figure 2: Revela’s architecture. With a specific attention map, the embeddings of in-batch attention $\{\mathrm{h}_{i}^{l}\}_{i=1}^{B}$ can attend to the self-attention $\{\mathrm{e}_{i}^{l}\}_{i=1}^{B}$.
As show in Fig. 2, in Revela, the NTP for sequence $i$ is conditioned not only on its own preceding context $x_{<l}^{i}$ but also on every other document in the batch, $\{D_{j}\}_{j\neq i}$. Specifically, we introduce a new attention mechanism in the transformer blocks, i.e., in-batch attention (Sec. 3.2), which is supported by a specific mask design (Sec. 3.4). The attentions are weighted by the inter-sequence similarity computed by the retriever parameterized by $\Theta$, i.e., $\mathrm{Sim}(D_{i},D_{j})$ (Sec. 3.3), where $\sum_{j\neq i}^{B}\mathrm{Sim}(D_{i},D_{j})=1$. The retriever is dynamically optimized as the similarity is updated jointly with the NTP objective
@@ 3.2 In-batch Attention
Revela augments the transformer block with an in-batch attention mechanism to incorporate context from other sequences. We denote the output of $D_{i}$ at the $l$-th layer as $[\mathrm{e}_{i}^{l};\mathrm{h}_{i}^{l}]$, where $\mathrm{e}_{i}^{l},\mathrm{h}_{i}^{l}\in\mathbb{R}^{L\times d}$ represent the outputs from self-attention and in-batch attention, respectively. For simplicity, we omit layer normalization and feed-forward layers, so the input to the two attention modules in the $l$-th layer is $[\mathrm{e}_{i}^{l-1};\mathrm{h}_{i}^{l-1}]$.
@@
Normalization by value norms prevents any single token from dominating the attention by scaling the output $\mathrm{b}^{l}_{ij}$ with the norm of $V_{j}^{e}$, thereby encouraging the cross-document attention to focus on sequence-level semantics izacard2023atlas . This operation, referred to as V-normalization, is computed as
where $\epsilon$ is a small constant for numerical stability.
Weighting by cross-document similarity aggregates the normalized attention outputs $\tilde{\mathrm{b}}^{l}_{ij}$ using the similarity scores $\mathrm{Sim}(D_{i},D_{j})$ computed by the retriever
@@ 3.4 Implementation Details
As described earlier, Revela adapts the classical transformers by additionally including in-batch attention, which builds upon standard self-attention, as shown in Eq. 5. For the minimum modifications to the existing transformer’s implementation, we take the computation of $\mathrm{e}$ and $\mathrm{h}$ as duplicating documents and adjusting the attention mask, as illustrated in Fig. 2. In this way, the embeddings $\{\mathrm{h}_{i}^{l}\}_{i=1}^{B}$ produced by in-batch attention can be obtained by applying full attention over the self-attention outputs $\{\mathrm{e}_{i}^{l}\}_{i=1}^{B}$ and aggregating them.
@@ 4.1 Experimental Setups
To comprehensively evaluate our proposed framework, we benchmark Revela on two diverse datasets: BEIR (thakur2021beir, ), a heterogeneous benchmark covering multiple domains for general information retrieval, and CoIR (li2024coir, ), a comprehensive benchmark designed for code-specific retrieval tasks. A more detailed introduction is listed in Appendix A.1.
@@
We use two different recipes of raw texts for the general and code-related retrieval mentioned above. The examples are presented in Appendix A.2. In this section, where we study the data independently, we also study the impact of the mixture of these two datasets in Sec. 5.2.
  • BEIR: We segment the passages in the Wikipedia corpus [^2] https://huggingface.co/datasets/Tevatron/wikipedia-nq-corpus into chunks of at most 120 words, each comprising complete sentences. Given a set of passages $\{d_{1},d_{2},\ldots,d_{n}\}$, where each passage $d_{i}$ is divided into a sequence of chunks $(d_{i1},d_{i2},\ldots,d_{im_{i}})$, the chunks are interleaved in the order $(d_{11},d_{21},\ldots,d_{n1},d_{12},d_{22},\ldots)$, and grouped sequentially into batches of size 16. In total, we sample 320,000 batches for training.
  • CoIR: Similarly, we chunk a set of code-related corpora (wang2024coderagbench, ), including the posts in the StackOverflow forum (weber2024redpajama, ), online tutorials (overwijk2022clueweb22, ), and library documentations (zhoudocprompting, ), for CoIR. The batch size is maintained as 16. Overall, there are 358,763 training batches.
@@
In Revela, we jointly train a retriever and an LM using self-supervised learning of LM, i.e., NTP. We adopt LLaMA-3.2-1B (grattafiori2024llama, ) as the LM. To ensure a fair comparison across diverse unsupervised baselines and retriever architectures, we adopt a range of LMs with parameter sizes from 0.1B to 1B as the backbone models for the retrievers: SmolLM2-135M (allal2025smollm2, ), Qwen2.5-0.5B (yang2024qwen2, ), and LLaMA-3.2-1B. We follow the approach used in RepLLaMA (ma2024fine, ). We append the <eos> token to each sentence and use its corresponding embedding as the sentence representation. Additionally, we prepend the prefixes "Query: " and "Passage: " to queries and passages, respectively. For more details about model checkpoints, please refer to Appendix A.4.
@@
We include representative self-supervised baselines: Contriever (DBLP:journals/tmlr/IzacardCHRBJG22, ) trains a BERT-based retriever via contrastive learning on unsupervised pairs; RetroMAE (xiao-etal-2022-retromae, ) reconstructs masked passages from compressed representations; LaPraDor (xu-etal-2022-laprador, ) combines latent-pair contrastive pretraining on C4 (JMLR:v21:20-074, ) with BM25-enhanced dense retrieval. REPLUG (shi-etal-2024-replug, ) distills supervision signals from frozen language models into retrievers, i.e., modeling chunk-chunk similarity within a batch by distilling the perplexity from an LM, using one chunk as prior context for the other. We select REPLUG [^3] For more details about REPLUG replication, please refer to Appendix A.3. as the main baseline for three reasons: (1) REPLUG consists of a retriever and a decoder-only LM, which is similar to Revela. Other approaches adopt different architectures—for instance, the encoder-decoder structure in Atlas (izacard2023atlas, )—which may divert attention from our primary focus on the joint training of retrievers and LMs. (2) Architectural similarity ensures that comparisons remain generalizable across different scales. (2) As reported by Shi et al. shi-etal-2024-replug , REPLUG outperforms most prior methods and can serve as a representative of previous research in this line of work.
When benchmarking CoIR, we also include several supervised retrievers, as well as API-based models such as OpenAI-Ada-002 (Ada-2) and Voyage-Code-002 (Voyage-2), following original setups (li2024coir, ). Supervised retrievers include GraphCodeBERT (GCBERT) (guographcodebert, ) and UniXcoder (guo-etal-2022-unixcoder, ), which are finetuned on code-related datasets. For more details about the Huggingface checkpoints, please refer to Appendix A.4.

### v2

@@ Abstract
Dense retrievers play a vital role in accessing external and specialized knowledge to augment language models (LMs). Training dense retrievers typically requires annotated query-document pairs, which are costly to create and scarce in specialized domains (e.g., code) or in complex settings (e.g., requiring reasoning). These practical challenges have sparked growing interest in self-supervised retriever learning. Since LMs are trained to capture token-level dependencies through a self-supervised learning objective (i.e., next token prediction), we can analogously cast retrieval as learning dependencies among chunks of tokens. This analogy naturally leads to the question: How can we adapt self‑supervised learning objectives in the spirit of language modeling to train retrievers?
To answer this question, we introduce Revela, a unified and scalable training framework for self-supervised retriever learning via language modeling. Revela models semantic dependencies among documents by conditioning next token prediction on local and cross-document context through an in-batch attention mechanism. This attention is weighted by retriever-computed similarity scores, enabling the retriever to be optimized as part of language modeling. We evaluate Revela on domain-specific (CoIR), reasoning-intensive (BRIGHT), and general-domain (BEIR) benchmarks across various retriever backbones. Without annotated or synthetic query-document pairs, Revela surpasses larger supervised models and proprietary APIs on CoIR and matches them on BRIGHT. It achieves BEIR’s unsupervised SoTA with ~ 1000x less training data and 10x less compute. Performance increases with batch size and model size, highlighting Revela’s scalability and its promise for self‑supervised retriever learning.
00footnotetext: https://github.com/TRUMANCFY/Revela00footnotetext: [Uncaptioned image] https://huggingface.co/trumancai/Revela-3b
@@ 1 Introduction
Central to information retrieval are dense retrievers (Reimers & Gurevych, 2019; Karpukhin et al., 2020; Ma et al., 2024), which map queries and documents into high-dimensional vector spaces and determine relevance through similarity calculations. Typically, these models rely on carefully annotated query-document pairs and hard negatives for training. However, creating such high-quality training data requires substantial human annotation, which is labor-intensive and difficult to scale in complex, domain-specific scenarios such as law (Feng et al., 2024) and programming (Jimenez et al., 2024). This limitation stimulates the interest within the community to explore self-supervised approaches for training retrievers directly from unannotated raw texts, i.e., self-supervised retriever learning (Izacard et al., 2022).
@@
Modern LMs (Grattafiori et al., 2024), a successful case of self-supervised learning, are typically pretrained with the next-token prediction (NTP) paradigm, modeling dependencies among tokens within a single sequence. Analogously, retriever aims to model relationships among larger units—chunks of tokens—capturing more macroscopic dependencies. This motivates us to raise a key question: How can we learn a retriever within a self-supervised learning framework of language modeling? While NTP implicitly identifies the most relevant parts in the context during generation, conditioning language modeling of one sequence on others can serve as an effective proxy for modeling inter-sequence relationships. This can offer a novel and principled approach to self-supervised retriever learning.
@@
In this work, we introduce a self-supervised retriever learning paradigm – Dense Retriever Learning via Language Modeling, abbreviated as Revela. As illustrated in Fig. 1, Revela trains retrievers by simultaneously optimizing retrievers and LMs; different from conventional NTP, Revela learns the probability of a token given both the prefix in this sequence and all other sequences in the batch. Specifically, in addition to classical self-attention which restricts NTP to individual sequences, an in-batch attention mechanism enables sequences to attend to their neighbors in the same batch during language modeling. In this process, the retriever provides inter-sequence dependencies that modulate the in-batch attention weights, allowing it to be optimized with the LM during training. We split raw texts into chunks within each document and put these chunks into the same batch, motivated by the idea of hard negative samples in contrastive learning (Xiong et al., 2020).
@@
Figure 1: The framework of Revela. The retriever’s in-batch similarity scores are used as in-batch attention weights inside transformer blocks. The retriever is trained by optimizing the language modeling objective, i.e., NTP. The related patterns in red and purple sequences are highlighted in bold and underline. An example of training dynamics is illustrated at App. A.
We comprehensively demonstrate the effectiveness of Revela across three benchmarks in § 4: CoIR (Li et al., 2024), a benchmark tailored for code retrieval, BRIGHT (Hongjin et al., 2025), a reasoning-intensive benchmark across diverse domains, and BEIR (Thakur et al., 2021), a heterogeneous benchmark covering general domains. To do this, we train Revela on Wikipedia for general retrieval and on code-related corpora (Wang et al., 2025) for code retrieval, using pretrained transformers ranging from 135M to 8B parameters as retriever backbones paired with an LM. On CoIR, Revela outperforms a strong, 7B-parameter supervised retriever (E5-Mistral-7b-Instruct) by 2.8 points and surpasses the weakly-supervised baseline E5-PT (Wang et al., 2022) by 9.7 points at a similar scale. This is particularly noteworthy as both baselines was pre-trained on massive query-document pairs that encompass Revela’s training data. Furthermore, Revela matches proprietary models on BRIGHT and achieves parity with weakly-supervised E5 model on BEIR while using approximately 1,000x less training data and 10x less compute. These results establish Revela as a highly effective and efficient self-supervised solution. In § 5, we also show that scaling Revela via on larger retriever backbones, larger LMs, and larger batch size can yield greater gain over baselines. Compared to traditional contrastive learning, Revela exhibits stronger cross-domain generalization. Moreover, the mixed-data training allows the model to scale across multiple domains while guaranteeing high domain-specific performance. Collectively, the evidence highlights Revela’s robust scaling behavior and strong generalization across models, data, and domains. To this end, we summarize our contributions as follows:
@@
  • Without query-document pairs, Revela surpasses E5-Mistral-7b-Instruct on CoIR by 2.8% and outperforms unsupervised baselines by 9.7% at the comparable scale, while also matching the performance of proprietary APIs on BRIGHT, a challenging, reasoning-intensive benchmark.
  • Revela exhibits robust scalability across larger models, batch sizes, and mixed-domain data; it also exhibits stronger cross-domain generalization than unsupervised contrastive methods.
@@ 2 Related Works
Some methods leverage weak supervision from document corpora. Contriever (Izacard et al., 2022) applies contrastive learning, using passages from the same document as positives and in-batch examples as negatives. E5 (Wang et al., 2022) is trained on a massive dataset of query-document pairs from numerous sources. The primary drawback of this direction is the risk of overfitting to structural biases present in the training data. Other training strategies include distillation from existing retrievers and autoencoding. Distillation is exemplified by LaPraDoR (Xu et al., 2022), which enhances dense retrieval by incorporating signals from BM25 (Robertson et al., 2009). Autoencoding methods, such as RetroMAE (Xiao et al., 2022), learn embeddings via sentence reconstruction. A key drawback of autoencoding is the lack of pairwise supervision, which can cause overfitting to low-level details (Steck, 2020).
Our approach departs from the conventional query-document framework. Drawing inspiration from NTP in LMs, our method adapts the language modeling objective, shifting its focus from predicting adjacent tokens to capturing the inherent associations between entire texts (sequences of tokens).
@@
LM-driven query and document expansion, exemplified by Query2Doc (Wang et al., 2023), can be effective but is often computationally costly due to its need for powerful models. Augmenting LMs with relevant information from external knowledge stores not only improves performance in various NLP tasks but also enhances retriever learning. Atlas (Izacard et al., 2023) utilizes cross-attention scores between retrieved documents and the generated output as signal to train the retriever. However, Atlas uses an encoder-decoder architecture, which diverges from the prevailing trend of decoder-only models and requires costly periodical reindexing. In contrast, our work leverages a standard decoder-only architecture to model relationships between text chunks within and across documents, mitigating the need for reindexing.
With the rise of decoder-only LMs, REPLUG (Shi et al., 2024) enhances retrieval by prepending retrieved documents to the queries, training retrievers to produce query-document similarity aligned with the LM’s perplexity. However, the perplexity of frozen LMs is often poorly calibrated (Geng et al., 2024), resulting in suboptimal retriever learning. This issue can be optimized in Revela where retrievers and LMs are updated jointly during language modeling.
@@
In pre-training corpora, domain-specific knowledge is both scarce and rapidly evolving (Grossmann et al., 2023; Wen et al., 2025), making effective retrieval in specialized domains critically challenging. To enhance the adaptability of dense retrievers across domains, researchers have explored continual learning (Sachan et al., 2021; Oguz et al., 2022) and task-aware training (Cheng et al., 2023). However, these approaches still rely on query-document pairs from domain-specific datasets. Another approach seeks to simplify domain-specific retrieval for general-purpose dense retrievers. Cai et al. (2024) propose a divide-and-conquer strategy through a mixed-granularity retrieval framework, significantly enhancing dense retriever performance in scientific domains. Our work demonstrates Revela’s domain adaptation capability through language modeling on domain-specific raw texts.
@@ 3.1 Training Objectives
Figure 2: Revela’s architecture. With an attention map, the embeddings of in-batch attention $\{\mathrm{h}_{i}^{l}\}_{i=1}^{B}$ can attend to the self-attention $\{\mathrm{e}_{i}^{l}\}_{i=1}^{B}$.
As shown in Fig. 2, in Revela, the NTP for sequence $i$ is conditioned not only on its own preceding context $x_{<l}^{i}$ but also on every other document in the batch, $\{D_{j}\}_{j\neq i}$. Specifically, we introduce a new attention mechanism in the transformer blocks, i.e., in-batch attention (§ 3.2), which is supported by a specific mask design (§ 3.4). The attentions are weighted by the inter-sequence similarity computed by the retriever parameterized by $\Theta$, i.e., $\mathrm{Sim}(D_{i},D_{j})$ (§ 3.3), where $\sum_{j\neq i}^{B}\mathrm{Sim}(D_{i},D_{j})=1$. The retriever is dynamically optimized as the similarity is updated jointly with the NTP objective
@@ 3.2 In-batch Attention
Revela augments the transformer block with an in-batch attention mechanism to incorporate context from other sequences. We denote the output of $D_{i}$ at the $l$-th layer as $\mathrm{e}_{i}^{l};\mathrm{h}_{i}^{l}]$, where $\mathrm{e}_{i}^{l},\mathrm{h}_{i}^{l}\in\mathbb{R}^{L\times d}$ represent the outputs from self-attention and in-batch attention, respectively. For simplicity, we omit layer normalization and feed-forward layers, so the input to the two attention modules in the $l$-th layer is $\mathrm{e}_{i}^{l-1};\mathrm{h}_{i}^{l-1}]$.
@@
Weighting by cross-document similarity aggregates the attention outputs $\mathrm{b}^{l}_{ij}$ using the similarity scores $\mathrm{Sim}(D_{i},D_{j})$ computed by the retriever
$\mathrm{b}^{l}_{i}=\sum_{\begin{subarray}{c}j=1,j\neq i\end{subarray}}^{B}\mathrm{Sim}(D_{i},D_{j})\,\mathrm{b}^{l}_{ij}.$ (6)
@@ 3.4 Implementation Details
As described earlier, Revela adapts the classical transformers by additionally including in-batch attention, which builds upon standard self-attention, as shown in Eq. 5. For the minimum modifications to the existing transformer’s implementation, we take the computation of $\mathrm{e}$ and $\mathrm{h}$ as duplicating documents and adjusting the attention mask, as illustrated in Fig. 2. In this way, the embeddings $\{\mathrm{h}_{i}^{l}\}_{i=1}^{B}$ produced by in-batch attention can be obtained by applying full attention over the self-attention outputs $\{\mathrm{e}_{i}^{l}\}_{i=1}^{B}$ and aggregating them. Revela’s efficiency is discussed in § C.7.
@@ 4.1 Experimental Setups
To comprehensively evaluate our proposed framework, we benchmark Revela on three diverse datasets: CoIR (Li et al., 2024), a comprehensive benchmark designed for code-specific retrieval tasks, BRIGHT (Hongjin et al., 2025), a retrieval benchmark requiring intensive reasoning to retrieve relevant documents spanning diverse domains, and BEIR (Thakur et al., 2021), a heterogeneous benchmark covering multiple domains for general information retrieval. A more detailed introduction is listed in § B.1.
@@
One of the earlier baselines for weakly supervised retrievers is E5 (Wang et al., 2022), which collected 1.3B text pairs from diverse sources such as StackExchange, Wikipedia, Reddit, and scientific papers, and filtered them down to 270M pairs using handcrafted rules. For Revela, We simply convert two E5 pretraining subsets to our training corpus: StackOverflow for code-related retrieval (CoIR) and Wikipedia for reasoning-intensive and general retrieval (BRIGHT & BEIR). Data preparation and illustrative examples of these subsets are provided in § B.2.
  • CoIR: We segment a set of code-related corpora (Wang et al., 2025) into chunks of at most 120 words, each comprising complete sentences, including the posts in the StackOverflow forum (Weber et al., 2024), online tutorials (Overwijk et al., 2022), and library documentations (Zhou et al., 2023), for CoIR. The batch size is 16. Overall, there are 358,763 training batches.
  • BRIGHT & BEIR: Similarly, we segment the passages in the Wikipedia corpus [^1] https://huggingface.co/datasets/Tevatron/wikipedia-nq-corpus. Given a set of passages $\{d_{1},d_{2},\ldots,d_{n}\}$, where each passage $d_{i}$ is divided into chunks $(d_{i1},d_{i2},\ldots,d_{im_{i}})$, the chunks are interleaved in the order $(d_{11},d_{21},\ldots,d_{n1},d_{12},d_{22},\ldots)$ and then grouped sequentially into batches of size 16. In total, we sample 320,000 batches from 339,409 documents for training. Notably, a single batch may contain chunks from different documents, highlighting the flexibility of batch construction in Revela.
@@
Revela jointly trains a retriever and an LM using the NTP objective. We adopt LLaMA-3.2-1B (Grattafiori et al., 2024) as the LM. To ensure a fair comparison across diverse unsupervised baselines, we adopt a range of LMs with parameter sizes from 0.1B to 3B as the backbone models for the retrievers: SmolLM2-135M (Allal et al., 2025), Qwen2.5-0.5B (Yang et al., 2024), LLaMA-3.2-1B and LLaMA-3.2-3B. [^2] For computational efficiency, we set smaller batch size for the 3B and 8B model, e.g., 8. For BRIGHT, we specifically include LLaMA-3.1-8B as the retriever’s backbone for the comparison with large LM-based and proprietary retrievers. We follow the approach used in RepLLaMA (Ma et al., 2024), appending the <eos> token to each sentence and use its corresponding embedding as the sentence representation. Additionally, we prepend the prefixes ”Query: ” and ”Passage: ” to queries and passages, respectively. For more details about model checkpoints, please refer to § B.4.
@@
We include representative self-supervised baselines: E5-PTlarge (E5-PT; Wang et al. 2022) is trained with a contrastive objective, leveraging weak supervision signals from a curated large-scale dataset of text pairs (1.3B raw pairs, filtered to 270M) spanning multiple domains (e.g., code) and covering Revela’s training corpus. REPLUG (Shi et al., 2024) distills supervision from a frozen LM into a retriever by using LM perplexity to model within-batch chunk–chunk similarity, conditioning one chunk on the other. We adopt REPLUG [^3] For replication details, see § B.3. as our main baseline because (1) it matches Revela’s retriever as decoder-only LM design, unlike encoder–decoder systems such as Atlas (Izacard et al., 2023), keeping the focus on joint retriever–LM training; (2) this architectural match makes comparisons generalizable across scales; and (3) REPLUG outperforms most prior methods, making it representative of this line of work (Shi et al., 2024).
For CoIR, we also include several supervised retrievers, as well as API-based models such as OpenAI-Ada-002 (Ada-2; Neelakantan et al. 2022) and Voyage-Code-002 (Voyage-2; Voyage AI 2024), following the original setup (Li et al., 2024). Supervised retrievers include UniXcoder (UniX; Guo et al. 2022), which is finetuned on code-related datasets; BGE-M3 (BGE; Chen et al. 2024), a supervised model pretrained and finetuned on text pairs (including code); and E5-Mistral-7B-Instruct (E5-Mistral; Wang et al. 2024).
For BRIGHT, we use several strong baselines: API-based models from text-embedding-3-large (OpenAI), cohere-embed-english-v3.0 (Cohere; Cohere 2023), and voyage-large-2-instruct (VoyageAI), as well as E5-Mistral.
For BEIR, we include Contriever (Izacard et al., 2022), a BERT-based retriever trained via contrastive learning on unsupervised pairs, and LaPraDor (Xu et al., 2022) uses latent-pair contrastive pre-training on C4 (Raffel et al., 2020) and fuses dense scores with BM25 via lexicon-enhanced dense retrieval.
For more details about the Huggingface checkpoints, please refer to § B.4.